A flute player hears four beats per second when she compares her note to a 523 Hz tuning fork (the note C). She can match the frequency of the tuning fork by pulling out the "tuning joint" to lengthen her flute slightly.

What was her initial frequency?
Answer #1
Concepts and reason

The concept used to solve this problem is beat frequency.

Use the expression for beat frequency to find the initial frequency of the flute player.

Fundamentals

Beat can be defined as “the interference pattern between two sounds of slightly different frequencies”.

The expression for the beat frequency is,

fbeat=f1f2{f_{beat}} = {f_1} - {f_2}

Here, beat frequency is fbeat{f_{beat}}, final frequency is f2{f_2}, and initial frequency is f1{f_1}.

The expression for the beat frequency is,

fbeat=f1f2{f_{beat}} = {f_1} - {f_2}

Rearrange the above expression for f1{f_1}.

f1=fbeat+f2{f_1} = {f_{beat}} + {f_2}

The expression for the initial frequency is,

f1=fbeat+f2{f_1} = {f_{beat}} + {f_2}

Substitute 4beats/s4\,{\rm{beats/s}} for fbeat{f_{beat}} and 523Hz523\,{\rm{Hz}} for f2{f_2}.

f1=(4beats/s)+(523Hz)=527Hz\begin{array}{c}\\{f_1} = \left( {4\,{\rm{beats/s}}} \right) + \left( {523\,{\rm{Hz}}} \right)\\\\ = 527\,{\rm{Hz}}\\\end{array}

Therefore, the initial frequency is 527Hz527\,{\rm{Hz}}.

Ans:

The initial frequency of the flute player is 527Hz{\bf{527}}\,{\bf{Hz}}.